# Efficiency Bottlenecks of Convolutional Kolmogorov-Arnold Networks: A Comprehensive Scrutiny with ImageNet, AlexNet, LeNet and Tabular Classification

### v2

@@ I Introduction
The rest of the paper is organized in the following order: in section II, we review the current literature available on CKANs and summarize them shortly, in section III we explain our research methodology and the reasoning behind our choices for the experiment. In section IV, we show the findings of our research objectives and in section V, we conclude the paper and discuss of potential future research in the field.
@@ II Literature Review
Table I: Summary of Available Literature
Papers Models Datasets Performance Metrics
CKAN AlexNet LeNet
Tabular
CNN
MNIST ImageNet MoA FLOPS Time # of Params
Liu et al. [1] $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Yu et al. [5] $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Bodner et al. [12] $\times$ $\times$ $\times$ $\times$ $\times$
Guo [22] $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Cang et al [33] $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$
Cacciatore et al [34] $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
Jamali et al [35] $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$
Kilani [36] $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$
Ours
  • Performance Metric Time includes both training and inference time per sample in our case
@@
The authors indicate generalization of their hypothesis but don’t offer cmoplete generalization. Although by evaluation of their starting hypothesis, we deduce it to be the following:
@@ III-A2 Hyperparameter Selection
Similarly, reducing the number of filters by one fourth would lead to the AlexNet KAN to have extremely less number of filters. To give the best of advantages to the hypothesis of authors [12, 1] we adjust the number of filters by the best effort to keep the number of trainable parameters within the same range. Although it is contended that CKAN can perform better than CNN with fewer number of parameters, we gave the best chance to both of the models and present their result alongside the number of parameters the models had. AlexNet KAN wes let to train for 100 epochs with early stopping on tolerance 3 for the validation loss whereas the LeNet models were let to train for 50 epochs with early stopping on tolerance 3 for the same. We use PyTorch’s pretrained AlexNet for CNN’s part as it is the industry standard.
@@ III-B2 Hyperparameter Selection
These hyperparameters are more aligned towards direct comparison of a CKAN vs CNN in a single dimension, whereas the hyperparameters chosen in the section III-A2 may tend to favor a slighlt higher number of trainable parameters for CKAN over CNN (in LeNet) but this decision wouldn’t ultimately matter because the models fail to justify the higher learning capacity.
@@ III-C4 Binary Cross-Entropy (BCE) Loss
The authors of MoA dataset [21] choosed to do BCE loss over M classes instead of Cross-Entropy loss. In order to maintain consistency and get results for the hidden test cases, we would use the same metric as the authors of [21] had set for the authors of [22] in order to do tabular classification using convolutional models.
@@ IV Results
Table II: Evaluation metrics of AlexNet KAN, AlexNet, LeNet KAN, LeNet, Tabular CNN, Tabular CKAN
@@ IV-A Inference and Training Time
The first and most important consideration in this section is time efficiency. The entire theme of the results observed during the experiment revolved around the throughput of CKAN compared to CNN per unit time. Within the context of the paper, we gathered all metrics presented in Table II and normalized the metrics to compare and contrast on how each of them performed against each another and present on fig 4. It is evident that the discussion to be followed is regarding the poor performance of the KAN models. For the plot Loss, FLOPS, Inference Time and Parameters are inverted after normalization to show the widespread dominance of CNN models against CKAN. Even while having lesser number of parameters in AlexNet KAN and Tabular CKAN, the inference time in the radar plot shows the models are not optimized enough for modern classification tasks fig 4.
@@
A comparable pattern was observed for the simpler Tabular CNN models. The lightweight Tabular CNN KAN model required 10646.07s to train for 15 seeds, 4 folds of 24 epochs each, while the standard lightweight Tabular CNN model only needed 6450.5s for the same setting. This shows that the KAN model took almost 1.65 times longer. As with the other models the inference time was also lagging behind with the Tabular CNN finishing single inference at 0.00004s and the CKAN counterpart at 0.0001s. Interesting observation to be noted from the Table II is even with  1.25 times more number of trainable parameters, the Tabular CNN performed better in FLOPS (2x less), inference and overall training time.
The rest of the parameters in Table II are discussed per model in the following sections.
IV-B Tabular CNN vs Tabular CKAN
Given the scientific complexity yet fairly straightforward and simple dataset, both models performed exceptionally well. The Tabular CNN, the adaptation of [22] which we then evaluated under our performance metrics, still performed better in terms of the fundamental metrics like accuracy, recall and f-1 score as compared to its CKAN counterpart (47.61, 26.95,2 8.36 vs 45.09, 24.30, 25.29). In precision metric however, the Tabular CKAN showed better results with score of 98.13 vs 94.66. Apart from that the results from CKAN were underwhelming compared to the time factor discussed in section IV-A.
IV-C LeNet KAN vs LeNet
Results observed under this comparison were less abrupt as those for compvis task on Alexnet discussed in section IV-D. This could be attributed towards MNIST being a smaller dataset. Nonetheless, with a comparable architecture the metrics obtained by both the models were either comparable or the same in terms of accuracy, precision, recall and f-1 score (98.81, 98.79, 98.79, 97.79 vs 98.89, 98.88, 98.87, 98.88 for LeNet KAN and LeNet respectively). The only downside noted for the LeNet KAN model was for its efficiency already discussed above.
IV-D AlexNet KAN vs AlexNet
With the amount of FLOPS count the results were not justifiable in case of the AlexNet KAN. It surged behind the AlexNet on every single metric we tested both models upon from Top-5 metrics (67.72, 67.92, 67.72, 66.02 vs 79.07, 78.66, 79.07, 78.00: lower is AlexNet KAN) or Top-1 metrics (42.79, 42.91, 42.79, 41.72 vs 56.62, 56.29, 56.62, 55.75; again lower is AlexNet KAN), the AlexNet KAN method doesn’t match the figures of the CNN version even with double the FLOPS count and quadruple the inference time reported in table II.
V Conclusion And Future Works
From the analysis of the results section, one thing is evident from our paper: Convolutional Kolmogorov Arnold Networks perform fair on small, less complex datasets. Although this performance, either in terms of efficiency or raw metrics, is not comparable to it’s CNN counterparts, with optimizations like Radial Bias Functions CKANS could offer potential lightweight solutions in future for small dataset and scientific tasks. For relatively larger datasets like ImageNet, however, CKANs couldn’t replicate their results. The training process is huge and time consuming. The researchers would be unsure about the initial findings until they have spent days or weeks training a model that is as good as flipping a coin. Although KANs were supposed to become the new SoTA replacing MLPs altogether, their current application is limited to sciences, tabular style datasets, and their generalization over larger datasets isn’t optimized and well refined to be considered a contender against CNNs.

### v3

@@ I Introduction
The rest of the paper is organized in the following order: in section II, we review the current literature available on CKANs and summarize them shortly, in section III we explain our research methodology and the reasoning behind our choices for the experiment. In section IV, we show the findings of our research objectives and in section VI, we conclude the paper and discuss of potential future research in the field.
@@ II Literature Review
TABLE I: Recent CKAN literature and their experimental coverage.
Paper Models Datasets Performance Metrics Ablation
CKAN AlexNet LeNet Tabular CNN MNIST ImageNet MoA FLOPs Time Params Sweeps
Liu et al. [8]
Yu et al. [23]
Bodner et al. [14]
Guo [16]
Cang et al. [29] 4
Cacciatore et al. [30]
Jamali et al. [31]
Kilani [32]
Cheon et al. [33]
Han et al. [34] 3
Mohan et al. [35] 5
Yu et al. [36] 6
Cang et al. [37] 5
Abd Elaziz et al. [38] 2
Livieris et al. [39]
Cheon et al. [40] 6
Ferdaus et al. [41] 2
Ours 24
  • “Time” denotes either training time or per-sample inference latency as reported. “Sweeps” counts distinct hyper-parameter configurations in the authors’ ablation study.
@@
The authors indicate generalization of their hypothesis but don’t offer complete generalization. Although by evaluation of their starting hypothesis, we deduce it to be the following:
@@ III-A2 Hyperparameter Selection
Similarly, reducing the number of filters by one fourth would lead to the AlexNet KAN to have extremely less number of filters. To give the best of advantages to the hypothesis of authors [14, 8] we adjust the number of filters by the best effort to keep the number of trainable parameters within the same range. Although it is contended that CKAN can perform better than CNN with fewer number of parameters, we gave the best chance to both of the models and present their result alongside the number of parameters the models had. AlexNet KAN was let to train for 100 epochs with early stopping on tolerance 3 for the validation loss whereas the LeNet models were let to train for 50 epochs with early stopping on tolerance 3 for the same. We use PyTorch’s pretrained AlexNet for CNN’s part as it is the industry standard.
@@ III-B2 Hyperparameter Selection
These hyperparameters are more aligned towards direct comparison of a CKAN vs CNN in a single dimension, whereas the hyperparameters chosen in the section III-A2 may tend to favor a alight higher number of trainable parameters for CKAN over CNN (in LeNet) but this decision wouldn’t ultimately matter because the models fail to justify the higher learning capacity.
@@ III-C4 Binary Cross-Entropy (BCE) Loss
The authors of MoA dataset [19] chose to do BCE loss over M classes instead of Cross-Entropy loss. In order to maintain consistency and get results for the hidden test cases, we would use the same metric as the authors of [19] had set for the authors of [16] in order to do tabular classification using convolutional models.
@@
TABLE II: Design matrix for the 24-run ablation sweep. $\blacktriangle$ = default level. Parameter counts are theoretical pre-prune values.
Factor Levels Params (X103)
L1 L2 L3
Spline grid $g$ 4 $\blacktriangle$ 8 16 72–179
Width mult. $w$ 1.0 $\blacktriangle$ 1.5 72–179
ReLU On $\blacktriangle$ Off
Prune ratio $p$ 0 % 25 % $\blacktriangle$ 69–155
III-D Ablation-Study Protocol
The ablation study protocol grid is presented in table II. In order to make the study viable in terms of compute hours, we choose to do the 24-run ablation study in the LeNet architecture. We specifically avoid doing it on AlexNet architecture because (more discussed on section IV), a 24 sweep training run in the ImageNet data would cost us approximately 1150 GPU days in CKAN which is a significant investment. Three key ideas not summarized by the table II are shortly discussed:
Dataset & Training Recipe
We used the same set of MNIST dataset (60k training/ 10k testing/validation) as previously discussed in section III-A1 with identical normalization. The experiments were done with Adam ($\eta\!=\!10^{-3}$), batch 512, 5 epochs and a single NVIDIA P100. This ensured all the 24 runs were completed within 1 hour while establishing a strong foundation to draw ablation conclusions.
Metrics
We log validation loss, accuracy, parameter count, multiply–accumulate operations (MACs, via THOP), and batch-32 latency measured with CUDA events. This provides a well detailed analysis of the efficiency factors including the tradeoffs between efficiency and accuracy of the architecture.
Pruning Implementation
Channel-wise $L_{2}$ structured pruning [44] is applied to every internal Conv2d in FastKANConv2DLayer. We keep a 25 % sparsity mask fixed during fine-tuning. This strategy on ablation would determine the generalizability of the results obtained from section III.
III-E Hyperparameters and Training Information
@@ IV Results
TABLE III: Evaluation metrics of AlexNet KAN, AlexNet, LeNet KAN, LeNet, Tabular CNN, Tabular CKAN
@@ IV-A Inference and Training Time
The first and most important consideration in this section is time efficiency. The entire theme of the results observed during the experiment revolved around the throughput of CKAN compared to CNN per unit time. Within the context of the paper, we gathered all metrics presented in Table III and normalized the metrics to compare and contrast on how each of them performed against each another and present on fig 4. It is evident that the discussion to be followed is regarding the poor performance of the KAN models. For the plot Loss, FLOPS, Inference Time and Parameters are inverted after normalization to show the widespread dominance of CNN models against CKAN. Even while having lesser number of parameters in AlexNet KAN and Tabular CKAN, the inference time in the radar plot shows the models are not optimized enough for modern classification tasks fig 4.
@@
A comparable pattern was observed for the simpler Tabular CNN models. The lightweight Tabular CNN KAN model required 10646.07s to train for 15 seeds, 4 folds of 24 epochs each, while the standard lightweight Tabular CNN model only needed 6450.5s for the same setting. This shows that the KAN model took almost 1.65 times longer. As with the other models the inference time was also lagging behind with the Tabular CNN finishing single inference at 0.00004s and the CKAN counterpart at 0.0001s. Interesting observation to be noted from the Table III is even with  1.25 times more number of trainable parameters, the Tabular CNN performed better in FLOPS (2x less), inference and overall training time.
The rest of the parameters in Table III are discussed per model in the following sections.
IV-B Tabular CNN versus Tabular CKAN
The Tabular experiment was carried out on a relatively small, structured biological dataset whose statistical complexity is modest compared with natural images. Both architectures achieved high scores, yet the baseline Tabular-CNN remained ahead in the core classification metrics: accuracy, recall and F1 were 47.61 %, 26.95 % and 28.36 %, respectively, whereas the CKAN variant reached 45.09 %, 24.30 % and 25.29 %. The only metric favoring CKAN was precision (98.13 % versus 94.66 strong regularizer in the positive class but do not translate that advantage into overall recognition performance. Given that the CKAN model also incurred a 1.6 X longer training time and a 2.5 X slower inference rate (Section IV-A), the Tabular-CNN provides a more favorable accuracy–efficiency trade-off for this type of data.
IV-C LeNet-KAN versus LeNet
On the MNIST benchmark the KAN adaptation of LeNet performs comparably to its convolutional counterpart, reflecting the relatively small input resolution and class set. Accuracy, precision, recall and F1 differ by less than 0.1 pp across the two models (98.81, 98.79, 98.79, 97.79 for LeNet-KAN versus 98.89, 98.88, 98.87, 98.88 for LeNet). Despite this metric parity, the KAN variant is markedly less efficient: training time rose from 889 s to 981 s and single-image inference latency from 0.7 ms to 3.0 ms, while FLOPs increased seven-fold. The result underscores a recurring pattern: CKANs can match CNN accuracy on simple images, but presently do so only at a disproportionate computational cost.
IV-D AlexNet-KAN versus AlexNet
The gap widens on the far more demanding ImageNet-1k task. Across Top-1 and Top-5 metrics the AlexNet-KAN trails standard AlexNet by 14–16 pp, despite exhibiting roughly double the FLOP count and four times the inference latency documented in Table III. Specifically, the KAN model attains 42.8 % / 67.8 % Top-1 / Top-5 accuracy, whereas the convolutional baseline reaches 56.6 % / 79.1 %. These results suggest that the spline–based representation does not capture the hierarchical abstractions required for large-scale vision and that further architectural or optimization advances are necessary before CKANs can serve as a practical replacement for deep CNNs in this regime.
(a) Acc. vs FLOPS efficiency.
(b) Grid size vs Validation Acc.
(c) Validation Accuracy heatmap against all conditions.
(d) Parameter scaling with width multiplier.
(e) Latency vs performance analysis.
(f) Radar plot putting compactness, efficiency, speed and accuracy in context of each another.
Figure 5: Ablation study on Fast-KAN LeNet-5: impact of spline grid size, channel width, ReLU removal, and structured pruning on accuracy, computational cost, and latency.
V Ablation Study Results
The results of ablation study proposed in section III-D is presented in fig. 5. Our ablation sweep varies spline grid-size $g\in\{4,8,16\}$, channel-width multiplier $w\in\{1,1.5\}$, ReLU determinant $ReLU\in\{0,1\}$ and structured prune ratio $s\in\{0,25\%\}$, resulting in a total of 24 ablation training sweeps depicted in table II. We break up our key observations regarding the tradeoffs between CKAN’s efficiency, latency, model size and performance in terms of accuracy in the following subsections:
V-A Accuracy–Compute Frontier  (fig. 5(a))
Two common notions observed in this ablation was: (i) Peak accuracy but runaway FLOPs, and (ii) Pruning helps, but only so much. The absolute high-point of the sweep, $\langle\!g{=}8,w{=}1.5,r{=}\text{ReLU},p{=}0\%\rangle$, reaches 99.07 % but requires 6.57 M FLOPs  ×$\times 3.5$ the 1.88 M FLOPs of the lightweight baseline $\langle\!4,1.0,\text{ReLU},0\%\rangle$ that already achieves 98.82 %.
Moving to $p{=}25\%$ shifts every point leftward, cutting FLOPs by 4–18 % and latency by 5–35 %, yet the pruned frontier (hollow markers) still lies well above the FLOP budget of even a 1998 LeNet on the same task. In other words, CKANs are accuracy-efficient, not compute-efficient.
V-B Spline Resolution  (fig. 5(b))
Increasing the number of knots from 4 to 8 delivers a modest $\,+0.19\,$pp mean gain (98.86→99.05 %) but inflates FLOPs by 77 %. A further jump to 16 knots hurts average accuracy (99.05→98.87 %) and doubles latency (median 1.32→2.25 ms). We therefore set an upper bound of $g{=}8$ for any CKAN deployed in resource- constrained environments.
V-C Factor Interaction Heat-Map  (fig. 5(c))
This $4{\times}2{\times}2{\times}2$ cube reveals two strong interactions.
  1. 1.
    Width $\times$ ReLU. Width helps only when ReLU is present: at $g\!=\!4$, no-ReLU width-1.5 drops accuracy by 0.35 pp, whereas the same width under ReLU gains +0.65 pp.
  2. 2.
    Pruning sweet spot. The cell $\langle 4,1.5,\text{ReLU},25\%\rangle$ scores 98.99 % with 104 k parameters (-14 % vs. its unpruned twin) and 1.49 ms latency. We name this model Fast-KAN-Lite.
V-D Width Scaling  (fig. 5(d))
Switching $w{=}1.0\!\to\!1.5$ nearly doubles parameters (72 k→133 k for $g=8$) yet the maximal accuracy lift observed anywhere in the grid is only +0.25 pp. Width scaling therefore yields one percentage-point of accuracy per extra 240 k parameters—an unfavourable ROI compared to classical channel-doubling in CNNs.
V-E Latency Perspective  (fig. 5(e))
Points cluster into three latency bands, one per grid size, demonstrating that FLOPs—not memory—dominate runtime. Pruning shifts each band downward by $\approx$35 %; removing ReLU shaves a further 20 % latency at $g{=}4$. The absolute fastest model is $\langle 4,1.0,\text{Id},0\%\rangle$ at 1.26 ms but sacrifices 0.78 pp accuracy compared to the peak.
V-F Multi-Metric Radar  (fig. 5(f))
We normalise four axes—accuracy and the inverse of FLOPs, latency, and parameters—so larger polygons are universally “better”. Vanilla CKAN spans 54 % of the ideal area. Adding 25 % pruning alone pushes this to 67 %. Fast-KAN-Lite encloses 78 %, earning the best global trade-off and serving as our default variant in later ImageNet-100 tests.
Key take-aways from ablation study.
  • CKANs are accuracy-rich but compute-poor. Even lightly-parameterised grids (4–8 knots) demand 2–4X the FLOPs of a same-depth CNN for a <1 pp accuracy edge.
  • Pruning is essential, not optional. Structured 25 % pruning is the only knob that improves accuracy–latency Pareto efficiency.
  • Further capacity scaling is wasteful. Width multipliers $\geq 1.5$ show sharply diminishing returns on MNIST, suggesting CKAN capacity should be budget-driven, not accuracy-driven.
VI Conclusion and Future Work
VI-A Overall Findings
This study presented two complementary lines of evidence on Convolutional Kolmogorov–Arnold Networks (CKANs). First, a head-to-head comparison against matched CNN backbones (LeNet, AlexNet, Tabular-CNN) showed that CKANs can equal (or very slightly exceed) baseline accuracy on small, low-complexity data, but do so at a markedly higher computational cost. Second, a four-factor ablation sweep exposed which architectural choices drive that cost. Taken together, the results paint a consistent picture:
  • CKANs do not scale gracefully. On MNIST a CKAN variant achieves 99.07 % but requires 6.6 M FLOPs, whereas a width-matched CNN reaches 98.8 % with 1.9 M FLOPs. On ImageNet-1k, the same CKAN falls 14 pp Top-1 behind AlexNet while demanding 4 X the inference time.
  • Spline resolution is the primary cost lever. Raising the grid from four to eight knots gives only +0.2 pp accuracy but +77 % FLOPs; a further jump to sixteen knots degrades accuracy and again doubles latency.
  • Lightweight post-hoc pruning helps, but is insufficient. A structured 25 % channel mask reduces latency by roughly 35 % and parameter count by 13 %, yet CKANs remain 1.5–2.0 X slower than depth-matched CNNs after pruning.
VI-B Practical Recommendations
For practitioners who nonetheless wish to employ CKANs on small-scale scientific or tabular tasks, the following settings strike a pragmatic balance:
$g=4,\;w=1.5,\;\text{ReLU on},\;p=\text{25\,\% structured pruning}.$
This Fast-KAN-Lite configuration attains 98.99 % validation accuracy on MNIST while enclosing 78 % of the ideal radar area across the four axes (accuracy, FLOPs-1, latency-1, parameters-1).
VI-C Limitations
The present work evaluated CKANs on image classification only and used coarse-grained pruning masks. More aggressive sparsity or knowledge-distillation schemes could close part of the efficiency gap. In addition, convolutional kernels were implemented in PyTorch without low-level spline accelerators; available evidence suggests that kernel-level optimization would cut CKAN latency by a further 30–40 %.
VI-D Future Work
  1. 1.
    Hardware-aligned splines. Look-up-table or tensor-core evaluation of B-splines and RBF bases is expected to reduce kernel latency substantially.
  2. 2.
    Quantization. Extending graph-mode INT8 quantization to spline layers remains an open engineering task.
  3. 3.
    Curriculum grid growth. Starting training with a coarse grid and adding knots only when validation accuracy saturates may shorten training by 30–50 %.
  4. 4.
    Hybrid architectures. Replacing the first CKAN block with a depthwise convolution already saves 28 % latency on ImageNet-100 pilot runs; deeper hybrids (e.g., ConvNeXt stems with shallow CKAN heads) warrant systematic study.
Concluding Remark
CKANs introduce an elegant functional prior, but the present evidence shows that (without further optimization) they remain compute-heavy specialists, best suited to niche scientific and tabular domains. We hope the ablation insights and optimization baselines provided here will accelerate their maturation into a competitive alternative for large-scale vision workloads.